OpenStudy (anonymous):

Can you help me factor these? 2x2 + 13x + 15, 2x2 + 27x + 13

OpenStudy (hba):

i suppose u asked the same question lexi

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

2x2+13x+15 have solutions are -3/2 and -5 2x2+13x+15=2(x+3/2)(x+5)=(2x+3)(x+5)